OpenStudy (anonymous):

A 175 kg cart is pushed along level ground for 18 m, with a force of 430 N and then released. a) Determine how fast was the cart been traveling when it was released? b) Determine how far will the cart travel after been released?

OpenStudy (jusaquikie):

i'm not sure but this is what i get. (a) First we need the acceleration which we use NSL to get which is ma=fa (175a=430) or a=2.46 at the release acceleration is constant so we can use kinematics to find v at the time it was released. so firtst for t we get: Xf-Xi=V0t+(1/2*at^2) pluging in what we have 18-0+=0t+(1/2)(2.46)t^2 rewrite as 18/1.23 = t^2 = \[\sqrt{14.65}\] or 3.83seconds Now we use have a=2.46 m/s^2 t=3.83s x=18m (v=v0+at) which is (v=0+2.46*3.83) or v=9.4 at the time the cart was released. (b)i don't know how to figure tis out beacuse i would think that we need some form of friction or negative acceleration or it would go on forever. i could be wrong but this is where i would start. good luck.

OpenStudy (anonymous):

a=2.46 now for the 2nd part I would try this one. v^2=2a(delta X)